# StrucSum: Graph-Structured Reasoning for Long Document Extractive Summarization with LLMs

### v1

@@ 4.3 Summary Faithfulness Evaluation
Although extractive summaries are constructed from source sentences, prior work has shown that they can still suffer from factual inconsistencies, such as incorrect coreference and misleading sentence combinations Zhang et al. (2023b). To assess the faithfulness and overall quality of summaries under different prompting strategies, we evaluate GPT4o-mini on the ArXiv and PubMed datasets (Table 4). We report results using two factual consistency metrics, namely FactCC Kryściński et al. (2019) and SummaC Laban et al. (2022), as well as G-Eval Liu et al. (2023), which averages scores across coherence, consistency, fluency, and relevance.
As shown in Table 4, incorporating structural information consistently improves both faithfulness and quality metrics compared to vanilla prompting. Among all strategies, NAP achieves the highest scores across both datasets, demonstrating its strength in preserving factual accuracy and coherence. CAP also brings consistent gains, while CGM shows more variable performance: modestly improving G-Eval scores while offering substantial input efficiency benefits (Section 5.2).
Overall, these results suggest that structural prompting, particularly NAP, significantly improves the factual reliability and quality of LLM-generated extractive summaries. Injecting structural cues into prompts leads to more trustworthy summarization in zero-shot settings.
@@
Arxiv GPT4o-mini 38.44 43.64 4.08
+ NAP 57.67 53.37 4.23
+ CAP 45.26 44.46 4.14
+ CGM 36.19 38.22 4.13
PubMed GPT4o-mini 44.16 63.81 4.20
+ NAP 55.44 64.64 4.32
+ CAP 49.43 63.28 4.25
+ CGM 45.46 62.85 4.23
Table 4: Evaluation results on summary factual consistency and holistic quality for GPT4o-mini with different prompting strategies on the ArXiv and PubMed datasets. We report scores from two automated factual metrics (FactCC, SummaC) and one general-purpose LLM-based evaluator (G-Eval).
@@ 4.4 Human Evaluation
We conducted a human evaluation on the PubMed dataset using a 1–5 scale for Fluency, Informativeness, Faithfulness, and Overall quality. As shown in Table 5, all proposed strategies (NAP, CAP, CGM) outperform vanilla prompting, with NAP consistently achieving the highest scores across all criteria. The improvements in human-perceived Overall quality (e.g., NAP: 3.74 vs. Vanilla: 3.30) aligns with the gains observed in the holistic G-Eval metric (Table 4). Furthermore, the improved human-rated Faithfulness (e.g., NAP: 3.95 vs. Vanilla: 3.61) is consistent with gains shown by automatic faithfulness metrics like FactCC and SummaC (Table 4). These consistent findings from both human and automatic evaluations verify that structure-aware prompting produces summaries with higher quality and reliability.
@@ 5.2 Token Usage Analysis
Prompt token usage is a critical consideration for practical LLM applications Sahoo et al. (2025). Figure 4 shows the normalized prompt lengths for our structure-aware strategies—Neighbor-Aware Prompting (NAP), Centrality-Aware Prompting (CAP), and Centrality-Guided Masking (CGM)—relative to the vanilla prompting method. Encoding more structural context, NAP and CAP increase input length by approximately 50–60% and 15–18%, respectively. In contrast, CGM, reduces token usage by 40–50% while still showing performance gains over the baseline methods (as detailed in Sections 4.2 and 4.3). Therefore, CGM can be considered as an efficient option for scenarios constrained by input budget or latency.
@@ 6 Conclusion
This paper presents StrucSum, a prompting-based framework for extractive summarization that augments LLM prompting with structural signals obtained from sentence-level graphs. By using three complementary strategies: NAP, CAP and CGM, StrucSum enables LLMs to reason more effectively over document structure in a zero-shot setting. Experiments across multiple benchmarks show that structure-aware prompting substantially improves both summary quality and factual consistency, with NAP achieving the most consistent gains. Overall, this study shows that structure-aware prompting is an effective and scalable approach for enhancing LLM-based summarization. We hope that the findings from this work can inspire further exploration of structure-guided reasoning in summarization and related tasks.
@@
Appendix A Prompts
In this section, we provide representative prompt examples for the four main methods evaluated in this paper, including vanilla prompting and the three structure-aware strategies: NAP, CAP, and CGM.
@@ Appendix A Prompts
You are an expert in extractive summarization. Your task is to select the most important sentences from the document. On average, select around 4 key sentences. Sentence 1: "Text_S1" Sentence 2: "Text_S2" Sentence 3: "Text_S3" Sentence 4: "Text_S4" Return a list of selected sentence indices in JSON, for example: {"selected_sentences": [1, 3, 5]}
@@
You are an expert in extractive summarization. Your task is to select the most important sentences from the document. Use information about neighboring sentences to help reason about local context. On average, select around 4 key sentences. Sentence 1: "Text_S1" - 1-hop Neighbors: Sentence 2, 4, 7, 10… Sentence 2: "Text_S2" - 1-hop Neighbors: Sentence 1, 3, 6, 11… Sentence 3: "Text_S3" - 1-hop Neighbors: Sentence 2, 5, 9, 13… Return a list of selected sentence indices in JSON, for example: {"selected_sentences": [1, 3, 5]}
@@
You are an expert in extractive summarization. Your task is to select the most important sentences from the document. Use the centrality scores as a guide for identifying global important sentences. On average, select around 4 key sentences. Sentence 1 (Centrality: 0.68): "Text_S1" Sentence 2 (Centrality: 0.45): "Text_S2" Sentence 3 (Centrality: 0.72): "Text_S3" Return a list of selected sentence indices in JSON, for example: {"selected_sentences": [1, 3, 5]}
@@
You are an expert in extractive summarization. Your task is to select the most important sentences from the document. The document has been pre-filtered to contain only structurally salient sentences based on their graph centrality. On average, select around 4 key sentences. Sentence 1: "Text_S1" Sentence 4: "Text_S4" Sentence 7: "Text_S7" Sentence 9: "Text_S9" [ Masked Sentences: 2, 3, 5, 6, 8, 10, … ] Return a list of selected sentence indices in JSON, for example: {"selected_sentences": [1, 3, 5]}

### v2

@@ 4.4 Summary Faithfulness Evaluation
Extractive summaries can still contain factual inconsistencies, such as coreference errors or misleading sentence combinations Zhang et al. (2023b). We evaluate summaries generated by GPT4o-mini on ArXiv and PubMed (Table 4) using three metrics: FactCC and SummaC for factual consistency, and G-Eval for summary evaluation with LLM-as-a-judge.
Across both datasets, all structure-aware strategies improve over vanilla prompting. NAP consistently achieves the highest scores on FactCC and SummaC, indicating its strong effectiveness in maintaining factual accuracy. It also slightly outperforms other methods on G-Eval. CAP and CGM show more metric-specific strengths: CAP performs competitively on G-Eval and FactCC, while CGM achieves strong gains on SummaC, particularly on PubMed.
We infer that NAP mainly reduces local reasoning errors rather than increasing reference overlap. By exposing each sentence’s 1-hop neighbors, NAP supplies a short-range discourse context that constrains the LLM’s sentence stitching. This directly lowers contradiction and attribution mistakes, which FactCC and SummaC are designed to detect, thereby yielding large gains on faithfulness-oriented metrics.
These results suggest that structural prompting enhances the faithfulness of extractive summaries. Among the strategies, NAP is the most effective overall, while CAP and CGM offer complementary benefits depending on the focus of the evaluation.
@@
ArXiv GPT4o-mini 38.44 48.64 4.08
+ NAP 57.67 56.61 4.23
+ CAP 45.26 49.07 4.14
+ CGM 36.19 55.72 4.13
PubMed GPT4o-mini 44.16 61.89 4.20
+ NAP 55.44 66.67 4.32
+ CAP 49.43 65.56 4.25
+ CGM 45.46 64.94 4.23
Table 4: Evaluation results on summary factual consistency and holistic quality for GPT4o-mini with different prompting strategies. We report scores from two automated factual metrics (FactCC, SummaC) and one general-purpose LLM-based evaluator (G-Eval).
NAP CAP CGM ROUGE SummaC
0.2781 0.5527
0.3005 0.6164
0.2998 0.5732
0.3034 0.6033
0.2974 0.5893
0.2979 0.5890
0.3002 0.6160
0.3002 0.5908
Table 5: Ablation and joint analysis of structure-aware strategies on the average of PubMed and ArXiv using GPT4o-mini. ROUGE (average of ROUGE-1/2/L) and SummaC are reported.
4.5 Ablation Study of Structural Strategies
We evaluate all valid combinations of NAP, CAP, and CGM on the average of PubMed and ArXiv using GPT-4o-mini. As shown in Table 5, individual strategies generally perform better than combined ones. CGM alone achieves the highest ROUGE score, while NAP alone gives the highest SummaC score. This suggests that CGM is more helpful for improving content coverage, whereas NAP is more effective for consistency. The full combination of NAP, CAP, and CGM performs in the middle for both metrics. In most cases, combining strategies does not lead to further improvement. These results indicate that each structural strategy focuses on different aspects of the summarization task, and using them together does not necessarily lead to better overall performance.
@@ 4.6 Human Evaluation
We conducted a human evaluation on the PubMed dataset using a 1–5 scale for Fluency, Informativeness, Faithfulness, and Overall quality. As shown in Table 6, all proposed strategies (NAP, CAP, CGM) outperform vanilla prompting, with NAP consistently achieving the highest scores across all criteria. The improvements in human-perceived Overall quality (e.g., NAP: 3.74 vs. Vanilla: 3.30) aligns with the gains observed in the holistic G-Eval metric (Table 4). Furthermore, the improved human-rated Faithfulness (e.g., NAP: 3.95 vs. Vanilla: 3.61) is consistent with gains shown by automatic faithfulness metrics like FactCC and SummaC (Table 4). The guidelines for our human evaluators are provided in Appendix B.
@@ 5.2 Token Usage Analysis
Prompt token usage is a critical consideration for practical LLM applications Sahoo et al. (2025). Figure 4 shows the normalized prompt lengths for our structure-aware strategies: neighbor-aware prompting (NAP), centrality-aware prompting (CAP) and centrality-guided masking (CGM) relative to the vanilla prompting method. Encoding more structural context, NAP and CAP increase input length by approximately 50–60% and 15–18%, respectively. In contrast, CGM, reduces token usage by 40–50% while still showing performance gains over the baseline methods (as detailed in Sections 4.3 and 4.4). Therefore, CGM can be considered as an efficient option for scenarios constrained by input budget or latency.
Figure 5: Spearman correlation coefficients between sentence centrality scores and selection rank under different prompting strategies on GPT-4o, shown separately for ArXiv and PubMed. Asterisks indicate statistical significance ($p<0.05$).
5.3 Correlation Analysis
We assess whether LLMs follow structural cues by computing Spearman correlations between sentence centrality and selection ranks. As shown in Figure 5, the prompts that incorporate centrality (via CAP or combined strategies) produce significant positive correlations ($p<0.05$), indicating the alignment between selection behavior and importance based on graphs. In contrast, NAP and CGM alone show nonsignificant trends ($p>0.05$), suggesting a limited effect without explicit importance signals. Although CAP successfully guides LLMs to prioritize central sentences, the results of Table 5 show that CAP-based configurations achieve smaller gains in ROUGE and SummaC compared to NAP and CGM. We infer that this may be because overemphasizing centrality limits the model’s flexibility to consider other informative sentences, which can affect both content coverage and consistency.
@@ 5.4 Structure-Only Prompt Format Analysis
Dataset $\boldsymbol{\theta}$ Avg #Nodes Avg #Edges Density
ArXiv 0.4 200 7792.6 0.1712
0.5 200 3197.6 0.0743
0.6 200 1076.5 0.0260
0.7 200 304.2 0.0072
0.8 200 85.6 0.0018
0.9 200 36.6 0.0007
PubMed 0.4 83 2700.5 0.3011
0.5 83 1410.3 0.1627
0.6 83 607.5 0.0744
0.7 83 210.3 0.0274
0.8 83 62.6 0.0078
0.9 83 26.9 0.0027
Table 8: Graph sparsity statistics averaged over 200 documents.
Graph Sparsity Analysis.
We analyze the sparsity of the constructed similarity graphs for the ArXiv and PubMed datasets under different similarity thresholds $\theta\in\{0.4,0.5,0.6,0.7,0.8,0.9\}$. Following standard definitions in network analysis Newman (2010), edge density is defined as |E|/(|V|(|V|1))|E|/(|V|(|V|-1)). As shown in Table 8, the resulting graphs are consistently sparse, and their densities decrease rapidly as the threshold increases (e.g., ArXiv: 0.1712 $\rightarrow$ 0.0007; PubMed: 0.3011 $\rightarrow$ 0.0027). Our experimental results confirm that the similarity graphs are sparse.
@@ 6 Conclusion
We propose StrucSum, a prompting-based framework for long-document extractive summarization that augments LLMs with structural signals from sentence-level graphs. StrucSum applies three strategies: namely, NAP, CAP, and CGM, to guide LLMs in a zero-shot setting. Experiments on three benchmarks show that structure-aware prompting improves both summary quality and factual consistency. StrucSum demonstrates that integrating simple graph priors into prompts enhances LLM reasoning over long documents. Future work may extend this framework to multilingual and dynamic graph reasoning.
Acknowledgment
This work was supported in part by National Science Foundation awards 2201428, 2004014, and 2138259, as well as by computing resources provided through the NSF ACCESS projects CIS250940 and CIS24061. We thank the anonymous reviewers for their valuable suggestions.
@@ Acknowledgment
Appendix A Prompt Examples
In this section, we provide prompt examples for the four main methods evaluated in this paper: Vanilla Prompting and the three structure-aware strategies: NAP, CAP, and CGM.
[…]
@@ Appendix A Prompt Examples
[…]
@@
[…]
@@
[…]
@@
Appendix B Human Evaluation Guidelines
Our human evaluation used two PhD students, each independently rating summaries using the rubric on a 1–5 scale for Fluency, Informativeness, Faithfulness, and Overall Quality. We report the average of their scores across examples. Here we list the full human evaluation guidelines used for our manual annotation study. including rating scales and definitions.
Score Fluency Informativeness Faithfulness Overall Quality
5 Grammatical, natural, and well-structured; reads smoothly with appropriate register; no typos or awkward phrasing. Covers all salient points and key details from the source; concise with minimal redundancy; no important omissions. Fully faithful to the source; no hallucinations, distortions, or unsupported inferences; entity, number, and event details are correct. Publication-quality summary: accurate, concise, coherent, and follows instructions; no substantive issues.
4 Mostly fluent with only minor issues (rare typos or slightly awkward phrasing) that do not impede comprehension. Covers most key points, with at most minor omissions or inclusion of unimportant details. Largely faithful; at most very minor imprecision that does not change meaning or introduce new facts. High-quality summary with minor issues in at most one dimension; clearly useful as-is.
3 Generally readable but with noticeable issues (grammar/style/word choice) that occasionally hinder the flow. Conveys several main points but misses multiple important details or includes some irrelevant/redundant content. Some unsupported statements, paraphrase drift, or partial contradictions; localized factual issues are present. Adequate but requires edits; there is clear room for improvement in two or more dimensions.
2 Frequent grammatical errors, unclear or choppy sentences; difficult to follow at times. Captures few key points; omits many important details and/or contains substantial irrelevant content. Multiple factual errors, contradictions, or hallucinated content; the meaning often diverges from the source. Low-quality summary with substantial issues; not reliable without significant revision.
1 Severely ungrammatical or incoherent; hard to understand. Fails to convey the main points; largely irrelevant or off-topic. Unfaithful to the source, with pervasive hallucinations or contradictions. Unusable summary that does not meet task requirements.
Table 9: Human evaluation rubric (1–5 scale) for Fluency, Informativeness, Faithfulness, and Overall Quality, used to assess GPT-4o-mini summaries under different prompting strategies.